Bob can load his truck in 24 min. It takes his brother John 40 min to load the truck. How long it take them to do the job together?


1
Expert's answer
2021-01-17T18:07:53-0500

1/24 - Bob load his truck in 1 min

1/40- John load the truck in 1 min 

1/24+1/40= 5/120+3/120=8/120=1/15 - Bob and John load the truck together in 1 min 


So it takes 15 minutes to do the job together.


Answer: 15 minutes